the dosage of medicine ordered by a doctor is 30 milliliters of a 12% solution. A nurse has available both a 30% solution and a 6% solution of this medicine. How many milliliters of each could be mixed to prepare this 30 ml dosage?

Answer by stanbon(75887)

the dosage of medicine ordered by a doctor is 30 milliliters of a 12% solution. A nurse has available both a 30% solution and a 6% solution of this medicine. How many milliliters of each could be mixed to prepare this 30 ml dosage?
--------
Equation:
0.30x + 0.06(30-x) = 0.12*30
-------------------------------------
30x + 6*30 - 6x = 12*30
24x = 6*30
x = (1/4)30
-----
x = 7.5 ml (amt of 30% solution needed)
------
30-x = 22.5 ml (amt. of 6% solution needed)
=============
Cheers,
Stan H.
=============